If household saving decreases by $4 million, business saving increases by $4 million, and the government budget deficit decreases by $4 million, then private saving ______ and public saving ______. Select one: a. increases; increases b. increases; decreases c. does not change; increases d. does not change; decreases..... expl why
Answer #1

Answer

c. does not change; increases

Budget deficit =G-T

and opposite

T-G =public saving

so

the deficit decreased means the public saving increased

=======

Private saving =sum of saving of households + sum of saving of businesses=-4+4=0

private saving does not change.